OpenStudy (anonymous):

x3/4 divided by x2/3

OpenStudy (anonymous):

hurry plz!!!

OpenStudy (irishboy123):

use this: a/b divided by c/d is the SAME as a/b multiplied by d/c

OpenStudy (anonymous):

??

OpenStudy (anonymous):

help

OpenStudy (ivettef365):

just subtract the exponents x(3/4 - 2/3)

OpenStudy (irishboy123):

ok a/b divided by c/d = a/b * d/c = ad/bc eg 1/4 divided by 3/7 = 1/4 multiplied by 7/3 = (1*7)/(4*3) = 7/12

OpenStudy (ivettef365):

3 2 9 - 8 1 -- - --- = --------- = ----- 4 3 12 12

OpenStudy (ivettef365):

so it would be x^1/12

OpenStudy (ivettef365):

Those are exponents right?

OpenStudy (irishboy123):

sorry, the question reads "x3/4 divided by x2/3" yet you have just (correctly!) subtracted them.

OpenStudy (anonymous):

thanks Irish boy

OpenStudy (ivettef365):

When you multiply exponents you have to add When you divide you subtract